OpenStudy (anonymous):

7-9) A 10 kg object (m1) initially moving to the right at 20 m/s makes a totally elastic head on collision with a 15 kg object (m2) which was initially moving to the left at 5 m/s. Find the final velocities of each object. HINT: Treat this as a one-dimensional problem.

OpenStudy (anonymous):

Answer is V1F: -10m/s and V2F: +15 m/s. Can someone explain this please? Thanks.

OpenStudy (kittiwitti1):

Define elastic? I would use the formula \[F = ma\]and then use the resulting info to find velocity. Also: \[a = \frac{ \Delta v }{ t } = \frac{ v }{ t }\]and this one: \[a = \frac{ F }{ m }\]

OpenStudy (kittiwitti1):

I borrowed the bottom two formulas from @Lessis so I get no credit for those. xD

OpenStudy (anonymous):

An elastic colission is a colission that conserves energy and momentum. So, that means that before and after the colission, the momentum and energy are equal.

OpenStudy (kittiwitti1):

Ah. Makes sense

OpenStudy (anonymous):

They bounce back like elastic balls loose no kinetic energy.

OpenStudy (anonymous):

lol, I'd help you if I knew how.

OpenStudy (kittiwitti1):

LOL. Thanks for the "want to help me" XDDDD

OpenStudy (anonymous):

Maybe once I have more knowledge on this subject, then I'll be able to help more people.

OpenStudy (kittiwitti1):

Well, of course. but I'm no help here what w/ the elastics, so @Lessis can help you more :)

OpenStudy (kittiwitti1):

See you on my other question XD

OpenStudy (anonymous):

Before the collision, the total momentum (The sum of the momentum of each particle) is the same as after. So: \[p _{1,1}+p _{2,1}= p _{1,2}+p _{2,2}\] You also have the same thing happening for the kinetic energy, since your potential energy is 0 (They're both on the ground). So: \[k _{1,1}+k _{2,1}= k _{1,2}+k _{2,2}\]

OpenStudy (kittiwitti1):

...why am I getting notifications on this....

OpenStudy (anonymous):

Notice that this only happens because your collisions are elastic, in inelastic collisions, some of your energy is lost (Due to heat, sound, etc).

OpenStudy (anonymous):

I think that makes sense. Thanks Lessis. @kittiwitti1 , once I close it, you won't be bothered anymore by this. Thank guys for you help. Appreciate it.

OpenStudy (anonymous):

Just remember, you have two velocities to get, and two equations to get them from, so you have to solve the system to get the answer. Good luck!